Calculating Departmental Overhead Rates Using Post-Allocation Costs

Valron Company has two support departments, Human Resources and General Factory, and two producing departments, Fabricating and Assembly.


Support Departments
Producing Departments
Human
Resources
General
Factory
Fabricating Assembly
Direct costs $160,000 $340,000 $114,600 $93,000
Normal activity:
Number of employees 60 80 170
Square footage 1,000 5,700 13,300

The costs of the Human Resources Department are allocated on the basis of number of employees, and the costs of General Factory are allocated on the basis of square footage. Valron Company uses the direct method of support department cost allocation. Solve for the allocated costs to Fabricating and Assembly using the direct method of support department cost allocation. The Fabricating Department overhead rate is based on normal activity of 82,000 machine hours. The Assembly Department overhead rate is based on normal activity of 160,000 direct labor hours.

    Job 316 required six machine hours in Fabricating and four direct labor hours in Assembly. Total direct materials cost $120, and total direct labor cost was $80.

Required:

1. Calculate the overhead rate for Fabricating based on machine hours and the overhead rate for Assembly based on direct labor hours. If required, round your answers to the nearest cent. Use the rounded values for subsequent calculations.

Overhead Rate
Fabricating department $ per mach. hr.
Assembly department $ per DLH

2. Using the overhead rates calculated in Requirement 1, calculate the cost of Job 316. If required, round your answer to the nearest cent.
$

3. What if Job 316 had required one machine hour in Fabricating and four direct labor hours in Assembly? Direct labor and direct materials costs remained the same. Calculate the new cost of Job 316. If required, round your answer to the nearest cent.
$
Answer #1
Support Departments Producing Departments
Human Resources General Factory Fabricating Assembly
Direct costs $160000 $340000 $114600 $93000
Allocation:
Human Resources (160000) - 51200 108800
General Factory - (340000) 102000 238000
Total after allocation $0 $0 $267800 $439800

Allocation of Human Resources Departments

Total number of employees of Producing departments= 80+170= 250

Fabricating= $160000*80/250= $51200

Assembly= $160000*170/250= $108800

Allocation of General Factory Departments

Total square footage of Producing departments= 5700+13300= 19000

Fabricating= $340000*5700/19000= $102000

Assembly= $340000*13300/19000= $238000

1) Fabricating department= Total cost allocated/Total machine hours

= $267800/82000= $3.27 per MH

Assembly department= Total cost allocated/Total labor hours

= $439800/160000= $2.75 per DLH

2) Overhead= (6*$3.27+4*$2.75)= $30.62

Cost of Job 316= Direct materials+Direct labor+Overhead

= $120+80+30.62= $230.62

3) Overhead= (1*$3.27+4*$2.75)= $14.27

Ne cost of Job 316= Direct materials+Direct labor+Overhead

= $120+80+14.27= $214.27